OpenStudy (vishweshshrimali5):

A simple algebra challenge \[\large{\color{red}{\text{THINK OUT OF THE BOX!!}}}\]

OpenStudy (vishweshshrimali5):

Let r and s be non zero integers. Prove that the equation \[\large{(r^2-s^2)x^2 - 4rsxy - (r^2-s^2)y^2 = 1}\] has no solutions in integers x and y

OpenStudy (vishweshshrimali5):

@Kainui @ganeshie8 @JungHyunRan @Miracrown @hartnn

OpenStudy (anonymous):

Would r and s represent trigonometric values?

OpenStudy (anonymous):

Or by simple algebra?

OpenStudy (vishweshshrimali5):

You have all the things given in the problem. You can use whatever you want but don't forget to justify your any assumption. I would prefer using simple algebra

OpenStudy (anonymous):

\[s = r \theta \]

OpenStudy (anonymous):

Alright thank you.

OpenStudy (vishweshshrimali5):

r and s are integers - remember this line :)

hartnn (hartnn):

b^2 -4ac = [16 r^2 s^2 +4 (r^2-s^2)^2 ] = 4[4 r^2 s^2 +r^4 -2r^2s^2 +s^4 ] = 4 [r^4 -2r^2s^2 +s^4] =4 (r^2-s^2)^2 which is a perfect square

hartnn (hartnn):

(4rs +/- 2(r^2-s^2))/2 (r^2-s^2) = -(r+/-s)^2 / (r^2-s^2) = -(r+s)/(r-s) or -(r-s)/(r+s)

hartnn (hartnn):

i guess thats thinking inside the box :P

OpenStudy (vishweshshrimali5):

Great but how will you prove the statement using this ?

OpenStudy (vishweshshrimali5):

Okay let me provide my solution...

OpenStudy (vishweshshrimali5):

We suppose that x and y are integers satisfying the given equation. Factoring the LHS, we get:

OpenStudy (vishweshshrimali5):

\[\large{((r-s)x - (r+s)y)((r+s)x + (r-s)y) = 1}\tag{1}\]

hartnn (hartnn):

wow! did anyone realize the blunder i made :P

OpenStudy (vishweshshrimali5):

Yeah +4 -2 = -2

OpenStudy (kainui):

I'm still trying to figure it out actually haha.

OpenStudy (anonymous):

I did this in my head just can't type it, my hands are broken

hartnn (hartnn):

lol , i took quadratic formula! applicable only for ax^2+bx+c = \(\huge 0\) not for ax^2+bx+c = 1 :P

OpenStudy (vishweshshrimali5):

Hahaha

OpenStudy (vishweshshrimali5):

Okay let me continue: as each factor in LHS of equation (1) is an integer, we can obtain the following: \[\large{(r-s)x - (r+s)y = \epsilon} \tag{2a}\] \[\large{(r+s)x + (r-s)y = \epsilon}\tag{2b}\] Where, \[\large{\epsilon = \pm 1}\]

OpenStudy (vishweshshrimali5):

Solving equations (2a) and (2b), I get: \[\large{x = \cfrac{r\epsilon}{r^2+s^2}}\] \[\large{y = \cfrac{-s\epsilon}{r^2+s^2}}\] Hence, we obtain : \[\large{(x^2+y^2)(r^2+s^2) = 1}\] So that: \[\large{r^2+s^2 =1}\] \[\large{\implies (r,s) = (\pm 1, 0)~or~(0, \pm 1)}\] which is not possible as r and s are both non zero. Thus, given equation has no integral solutions.

OpenStudy (vishweshshrimali5):

\[\huge{\color{blue}{\text{Q.E.D.}}~~\color{red}{\text{Vishwesh}}}\]

OpenStudy (vishweshshrimali5):

Well done @hartnn . Nicely tried :)

OpenStudy (vishweshshrimali5):

A mathematician who does not know to show off his work is an incomplete mathematician. That is why I used \(\large{\epsilon}\) in place of some other simple symbol like "k".